One-fifth of a number added to the number itself gives 42. Find the number. If I set this up as 1/5x + x = 42 and mutplied each side by 5, but this does not do it. I end up with 2x = 210. I know the answer is 35. How do I get to 6x = 210?
Answer by minerva(22)

well,u are adding incorrectly..
u see,
1/5x +x=42
multiply the eq by 5 thus.
5 X 1/5 x +5 X x=42x5
x+5x=210
6x=210
x=210/6
x=35.